OpenStudy (anonymous):

Choice #2: Describe each of the following properties of the graph of the Sine Function, f(0) = sin(0) and relate the property to the unit circle definition of sine. Amplitude Period Domain Range x-intercepts

OpenStudy (helder_edwin):

the period of the sine function is \(2\pi\) that is \[ \large \sin(x+2\pi)=\sin x \]

OpenStudy (helder_edwin):

u can get the sine of any number so its domain is \(\mathbb{R}\). for the range \(-1\leq\sin x\leq1\)

OpenStudy (helder_edwin):

for the x-intercepts \[ \large 0=\sin 0=\sin(k\pi) \] where k is any integer.